# Generated Data with Fake Privacy: Hidden Dangers of Fine-tuning Large Language Models on Generated Data

### v1

@@ 5 Related Work
5.1 Privacy Risks on Large Language Models
@@
The practical implications of these privacy risks have been well-documented in the literature. Many prior works have shown open-source LLMs leak significant parts of their training data. Various methods such as data deduplication and differential privacy [22] are proposed to mitigate the risks. However, these methods remain ineffective due to computational infeasibility of differentially-private stochastic gradient descent in addition to the evidence suggesting that memorization can still compromise privacy even in the absence of observable overfitting [32, 11].
These findings highlight the urgent need for further research into more realistic settings and the practical effectiveness of proposed mitigation techniques, and their impact on model utility. There remains significant uncertainty about whether fine-tuning exacerbates or mitigates memorization [11], as well as the broader effects of different training settings on privacy risks. Addressing these gaps is crucial for the development of more secure and privacy-preserving LLMs.
5.2 Privacy Risk with Synthetic or Generation Data
Using synthetic data in deep learning has been a common practice for numerous purposes [14, 37]. A prominent use case of synthetic, i.e. generated data is for training LLMs for downstream tasks [37, 34]. Recent works emphasized the efficacy of this use in terms of time and money [34]. Furthermore, the possibility of using LLMs locally for data generation appeared to be a remedy for concerns about privacy in multi-party computing settings [30].
However, the inherent risks of memorization and data leakage in LLMs raise concerns that fine-tuning on generated data may introduce significant, yet often overlooked privacy dangers. Specifically, generating synthetic data with a given prompt can lead to the reproduction of memorized data [33], a risk that parallels those seen in data extraction attacks [11]. Finetuning on the memorized data can further exacerbate the privacy risks, by leaking PIIs from the pretraining corpus of the target, or the generator models.
@@ 6 Conclusion
With the increment of data needed for fine-tuning, using the generated data for fine-tuning has become more popular these days. However, former research neglects the potential privacy risks when fine-tuning the model with the generated data. In this paper, we conduct experiments on two primary approaches to fine-tuning with generated data, supervised fine-tuning with unstructured generated data and self-instruct tuning, and evaluate the potential privacy risks when applying these fine-tuning pipelines. The results indicate LLMs can leak more private information on the related domain after fine-tuning with the generated data.
7 Ethics Considerations
This paper considers the potential privacy risks of the pre-training dataset for LLMs. It’s an important topic in LLMs Ethics. Apart from that, no ethical problems are related in this paper.
8 Compliance with the Open Science Policy
We will open-source the code after the acceptance.

### v2

@@ 5 Related Work
5.1 Privacy risks associated with LLMs
@@
Prior work has shown open-source LLMs leak significant parts of their training data. Various methods such as data deduplication and differential privacy [36] are proposed to mitigate the risks. However, these methods remain ineffective due to the computational infeasibility of implementing differentially private stochastic gradient descent. Additionally, evidence indicates that memorization can still compromise privacy, even in scenarios where observable overfitting is absent [49, 16].
These findings underscore the urgent need for further research into more realistic scenarios, the practical effectiveness of proposed mitigation techniques, and their potential impact on model utility. There remains significant uncertainty about whether fine-tuning exacerbates or mitigates memorization [16], as well as the broader effects of different training settings on privacy risks. We draw attention to this uncertainty and fill the important gap of fine-tuning on generated data, which is crucial for the development of more secure and privacy-preserving LLMs.
5.2 Privacy risks associated with synthetic data
Using synthetic data in deep learning has been a common practice for numerous purposes [20, 55]. A prominent use case of synthetic data is for training LLMs for downstream tasks [55, 52]. Recent works emphasized the efficacy of this use in terms of time and money[52]. Some works, like self-play fine-tuning [18], also demonstrate that using self-generated synthetic data can further improve the model’s performance. Furthermore, the possibility of using LLMs locally for data generation appeared to be a remedy for concerns about privacy in multi-party computing settings [47]. In practice, more and more developers adopting synthetic data for their fine-tuning, e.g., Llama-3 [2] and Tülu-3 [33], current state-of-the-art LLMs, both adopt LLM-generated data in their training data.
However, the inherent risks of memorization and data leakage in LLMs raise concerns that fine-tuning on generated data may introduce significant, yet often overlooked privacy dangers. Specifically, generating data with a given prompt can lead to the reproduction of memorized data [51], a risk that parallels those seen in data extraction attacks [16]. We note here that although has similar purposes, synthetic data is often created using statistical methods to mimic real data distribution. In contrast, LLM-generated data is produced from the model’s already-learned representations, potentially memorized patterns from pretraining data. This makes LLM-generated data potentially more susceptible to privacy attacks, as it can inadvertently amplify the leakage of the original pretraining data. Finetuning on the memorized data can further exacerbate the privacy risks, by leaking PIIs from the pretraining corpus of the target, or the generator models.
In our work, we investigated the membership inference risks of LLMs fine-tuned on domain-specific generated data, which had not been addressed before. We demonstrate our findings in the highly sensitive domain of law with prominent open-source models for research purposes.
@@ 6 Conclusion
With the growing data requirements for fine-tuning, the use of generated data has become increasingly common. However, previous research has overlooked the potential privacy risks associated with fine-tuning models using generated data. In this paper, we address this gap by conducting experiments on two primary fine-tuning approaches with generated data: supervised fine-tuning with unstructured generated data and self-instruct tuning. We then evaluate the potential privacy risks involved in these fine-tuning pipelines. The results indicate LLMs can leak more private information on the related domain after fine-tuning with the generated data.
Acknowledgements
We thank all anonymous reviewers for their constructive comments. This work is partially funded by the European Health and Digital Executive Agency (HADEA) within the project “Understanding the individual host response against Hepatitis D Virus to develop a personalized approach for the management of hepatitis D” (DSolve, grant agreement number 101057917) and the BMBF with the project “Repräsentative, synthetische Gesundheitsdaten mit starken Privatsphärengarantien” (PriSyn, 16KISAO29K).